ROTATIONAL SPECTRA AND INTERNAL ROTATION OF 1,1,1,2,2-PENTAFLUOROPROPRANE AND 2,2,2-TRIFLUOROETHYLMETHYLETHER

N. SAKAKIBARA, K. FUCHIGAMI, Y. TATAMITANI, \underline{T. OGATA}, Department of Chemistry, Faculty of Science, Shizuoka University, Shizuoka, Japan 422-8529.

The rotational spectra of CF_{3}CF_{2}CH_{3} and trans- and gauche-CF_{3}CH_{2}OCH_{3}, constitute alternative compounds to the CFCs, have been studied by microwave spectroscopy. The spectra were measured using Stark modulation and Fouriertransform microwave spectrometers at Shizuoka University. Both molecules exhibited the spectrum due to many excited states of the CF_{3} torsion. Accurate rotational constants and quartic centrifugal distortion constants have been obtained for the ground and torsional excited states up to v = 3. The barrier heights to internal rotation of the CF_{3} top have been determined for these molecules. Reasonable molecular structures have been derived from the observed rotational constants and were compared with those of ab initio calculations of GAUSSIAN MP2/6-311G** level.